OpenStudy (anonymous):

the lenght of a rectangle is 1 ft more than double the width, and the area of the rectangle is 28ft^2. find the dimensions of the rectangle.

OpenStudy (anonymous):

The formula for Area of rectangle is l*w It says that the length is 1+2w Width is still w the area is 28 Put L= 1+2w and W=w in the formula The equation will then be 28=(1+2w)*w Hope u can continue from here.If u have any questions u can ask.